4. Find the values of that satisfy the given equation: 11 -2 1 1 3 0 2-1 3 -11 1 -21
Answer #1

Consider 3 4 (1-0) [(2-1) 6-1) -83-0[-1-9]+(1) (9-363-] (1-d) (-21 +d2-3) - 3+3 (22-d) = 0 -21 +d²-3 + 2x = x3 +81-3 +6-3X=0